A coin collection totals $5.60, and consists of 35 coins. The number of dimes and the number of nickels is the same, and the number of quarters is one less than twice the number of dimes (or nickels). Find the number of each type of coin in the collection.
Answer by mananth(16946)

A coin collection totals $5.60, and consists of 35 coins. The number of dimes and the number of nickels is the same, and the number of quarters is one less than twice the number of dimes (or nickels).
...
A coin collection totals $5.60,
consists of 35 coins.
The number of dimes and the number of nickels is the same, =x
and the number of quarters is one less than twice the number of dimes 2x-1
..
x+x+2x-1=35
4x-1=35
4x=36
x= 9 dimes = nickels
quarters = 35-(9+9)
17 quarters
17*25+9*10+9*5= 560
,,,
m.ananth@hotmail.ca